# The physical basis of teleological evolutions

### v3

@@ Abstract
1 Foreword
The present work is about causal loops, of course a puzzling notion: in our natural way of seeing the world, we firmly believe that cause always precedes effect in time. In a causal loop, the effect, in turn, changes backward in time its previous cause. This may seem paradoxical; it is in fact the source of well-known paradoxes such as the grandfather one.
So, let us say right away that the causal loops we will be dealing with are quantum and of a special time-symmetric kind that should be imposed by quantum mechanics. Therefore, as long as the latter is consistent, they should be exempt from the paradoxes that plague the usual kind. They would be the physical reason for the quantum computational speedup, the fact that solving a problem in the quantum world can require fewer logical operations (computation steps) than in the classical world222 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was first introduced by Finkelstein $[1]$ in 1969. In 1982, Feynman $\left[2\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources. In 1985, Deutsch $\left[3\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup. . More in general, they would always be present in quantum correlations.
The present work is divided into two parts.
The first is a review and further clarification of the retrocausal approach to the quantum computational speedup $\left[4-8\right]$. It concerns a (time-symmetric) quantum causal loops description of quantum algorithms that provides a unified quantitative explanation of their quantum speedups. According to it, quantum algorithms are quantum superpositions of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. Then, she could use this knowledge to produce the solution with fewer computation steps. In equivalent terms, the computational complexity of the problem to be solved, or the dimension of the Hilbert space in which the search for the solution takes place, correspondingly reduces.
The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting from the advanced knowledge in question. We call this the advanced knowledge rule.
The first part of the work ends with the observation that, under their time-symmetric quantum causal loops description, quantum algorithms turn out to be teleological evolutions, namely evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution again) – a causal loop in fact.
@@ 1 Foreword
As is well known, the evolutions of the living were often thought to be teleological in character. The teleological notion was dismissed with the advent of modern science for the alleged absence of a physical basis. That was, of course, under the notion of causal evolution that is dominant in science also today and that the present work challenges.
The second part of the work is about the possibility that the time-symmetric quantum causal loops at the physical basis of the teleological character of quantum algorithms are also the physical basis of the teleological character of the natural evolutions for which the notion of teleological evolution was conceived. We will see that, under a quantum cosmological interpretation of the Anthropic Principle, this possibility turns out to be plausible.
@@ 2 Rationale
As the present work implies a fundamentally different way of seeing the role of causality in quantum correlation, we replace the usual introduction by a more extended rationale that describes this way of seeing in sufficient detail. Quantum correlation is of course the common feature of quantum algorithms (quantum correlation between the problem-setting and the corresponding solution), quantum nonlocality (quantum correlation between the outcomes of measuring two maximally entangled observables in a state of spatial separation of the respective subsystems), and a quantum cosmological interpretation of the Anthropic Principle (quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”).
The rationale is divided into two parts.
The first provides the concept of the time-symmetric quantum causal loops, showing that they stand at the physical basis of the quantum computational speedup and that, consequently, quantum algorithms respond to the notion of teleological evolution. This naturally raises the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
The second part of the rationale proposes an affirmative answer to this question. It is, of course, a speculation. It will imply seeing quantum features – the quantum speedup – in evolutions that can be the size of the Earth and last billions of years. We will show how this becomes possible under a quantum cosmological interpretation of the Anthropic Principle.
@@ 2.1 The time-symmetric quantum causal loops – the concept
We go now to the first part of the work. We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[9\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
We start with the simplest instance of the quantum algorithm. The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers; the number of the drawer in which to hide the ball – the problem-setting – is selected at random. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open from one to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if not even in that, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case, with the optimal quantum search algorithm of Grover, she always locates the ball by opening just one drawer, in a quantum superposition of the four possible drawer numbers. Always solving the problem by opening just one drawer would obviously be impossible in the classical case.
The fact that Alice always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
@@
We consider now $N$ drawers. We should use to Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty. Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance (in a mutually exclusive way) that the ball is in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
However, the above immediately raises a problem. The current quantum description of quantum algorithms, as that of any quantum process, is causal in character. In other words, causality always goes in the same time direction. Obviously, it is a description that cannot describe causal loops; we ascribe this to its incompleteness and show that it is completed by time-symmetrizing it.
First, we should emend a trivial incompleteness of the usual quantum description of quantum algorithms. It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is obviously incomplete because it lacks the initial measurement. Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings. In the case of Grover algorithm, where the solution is an invertible function of the problem-setting, we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement). Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
In view of going beyond Grover algorithm, we note that the one-to-one correlation between the two measurement outcomes is no restriction: if the solution is not an invertible function of the problem-setting, to have this correlation it suffices that, at the end, the problem-solver measures both the solution and the problem setting. The only constraint is that the quantum algorithm determines the solution with certainty.
We call the above quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops. It is the “less incomplete” quantum description that we should complete by the time-symmetrizing it.
In order that it generates the causal loops logically implied by the sheer existence of the quantum speedup, the time-symmetrization must be as follows.
@@
This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm. Note that, mathematically, i.e. disregarding the time-direction of causality, it is its ordinary quantum description back again.
Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure). As seen by the problem-solver (see Subsection 4.2.), from whom the outcome of the initial measurement should always be concealed (would give her information about the solution), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops.
As we can see, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered, but changes the structure of causality along it. As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. Then, she can use this knowledge to produce the solution with fewer computation steps. When she exploits it entirely, the quantum speedup is optimal.
@@
In the present work, we will always know that there is such a unitary evolution. It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[7\right]$. Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics. Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops. Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup, which is submitted to the same (classical) logic. In turn, quantum mechanics says that the only way of obtaining this is by time-symmetrizing the ordinary quantum description of the quantum algorithm. This, in turn, generates the description of the quantum superposition of causal loops in question. Therefore, as long as logic and quantum mechanics are consistent, these time-symmetric quantum causal loops should be exempt from the paradoxes that plague the usual kind.
@@
In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered. When the two measurements become simultaneous, the time-symmetrization in question must already be there to become an atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
@@
There is also the case that the unitary evolution does not build up any quantum correlation. It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation. Here maximal entanglement between the two observables is already there before the space-separation process. After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate. Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[8\right]$. These causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e. also physically) describe quantum nonlocality.

### v4

@@ Abstract
1 Introduction
@@
The notion of teleological evolution is that of an evolution toward a goal with an attractor in the same goal it will produce in the future. It is, of course, a puzzling notion since it necessarily implies a causal loop. As is well known, it was often applied to the evolutions of the living before the scientific revolution. With the advent of the latter, it was dismissed because of its alleged lack of physical basis. That was, of course, under the notion of causal evolution that is dominant in science even today.
We take the opposite position in the present work. We will show that, in quantum mechanics, there is a well defined physical basis to the notion of teleological evolution.
In the first place, we show that quantum algorithms are teleological evolutions, indeed evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
Of course, seeing this requires a quantum description of quantum algorithms different from the ordinary one, which is naturally that of a causal process. The latter quantum description, with causality always going in the same time-direction, cannot describe causal loops.
We show that the ordinary quantum description of quantum algorithms, although complete from the mathematical standpoint, is physically incomplete and is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by a quantum superposition of causal loops. In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps). Of course, we are dealing with an evolution that responds to the notion of teleological evolution: it is indeed toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again).
This completion of the quantum description provides a unified quantitative explanation of all the optimal quantum speedups achieved in solving oracle problems. Given an oracle problem, the number of computation steps needed to solve it in an optimal quantum way is that needed to solve it in a optimal logical way benefiting of the advanced knowledge of any of the possible halves of the information about the solution. Of course, there is no such an explanation in current quantum computing science. Obviously, we have a causal quantum description for each quantum algorithm, but there is no known way of unifying them.
This part of the work is a further development of the retrocausal approach to the quantum computational speedup $\left[1-5\right]$.
Of course, both the approach in question and the part of the work discussed so far are unconventional and potentially controversial. The notion of causal loop is puzzling and mostly extraneous to today’s physics. However, we believe that this time we have given a very simple and easily checkable demonstration that the quantum computational speedup is physically due to quantum causal loops.
Let us outline it. We consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer222 In the four drawers case, Grover algorithm yields the solution with certainty. In the following, by Grover algorithm we will always mean Long’s $\left[7\right]$ version of it, which does that always. . This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of viewing it.
It should already be clear from the above that that of the quantum speedup is a fundamental discovery that changes a long standing way of viewing things333 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[8]$ in 1969. In 1982, Feynman $\left[9\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources. In 1985, Deutsch $\left[10\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup. .
We can go now to our demonstration, which is surprisingly simple. Of course, we must adopt a way of thinking open to retrocausality.
The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers, in a mutually exclusive way one of the three pairs obtained by pairing the drawer with the ball with another drawer. In fact, if the ball is in the drawer opened, she has solved the problem. If not, it must be in the other drawer and she has solved the problem of locating the ball as well, having opened just one drawer.
Then the question becomes: from where does Alice always get the information that the ball is in a given pair of drawers? Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is obviously a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution by opening correspondingly fewer drawers. This, in the mutually exclusive $or$ of all the possible ways of taking half of the information about the solution.
We would argue that, with the above elementary logic, we have, so to speak, broken the sound barrier: the quantum speedup does logically imply causal loops. Then, one gos from logic to physics as follows.
Involving only the measurements of commuting observables (the problem-setting and the corresponding solution), quantum algorithms are submitted to the same (classical) logic. Therefore they must host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of an optimal quantum speedup; of course, logical mutual exclusivity of possibilities, moving to the quantum world, becomes their quantum superposition. We have already seen that the quantum description of this quantum superposition is generated by time-symmetrizing the ordinary quantum description of the quantum algorithm. This time-symmetrization must therefore be performed because, as we will see, is the only way of generating the causal loops that must be hosted in the quantum algorithm.
Let us add a few comments to the above.
By ordinary quantum description of quantum algorithms, we do not mean their usual quantum description. The latter is limited to the unitary process that solves the problem, where the quantum speedup resides, and the final measurement of the solution. It is trivially incomplete since it lacks the initial measurement. Since we need to complete the quantum description, in the first place we should emend this trivial incompleteness. We do this by extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
In the case of Grover algorithm, we obtain a unitary evolution between an initial and a final one-to-one correlated measurement outcomes encoding respectively the problem-setting (the number of the drawer with the ball selected by the initial measurement) and the corresponding solution of the problem. We have called it the ordinary quantum description of the quantum algorithm, ordinary since it is still a causal description that cannot describe causal loops. Note that any quantum algorithm that produces the solution with certainty can be put in this form. If the solution is not an invertible function of the problem-setting, it suffices that Alice, at the end, measures both the solution and the problem setting. This is a trivial but important step: of course, the usual quantum description of quantum algorithms would completely obscure the approach we are following.
Let us eventually note that we are dealing with the same kind of incompleteness of the ordinary quantum description that was ascribed by Einstein, Podolsky and Rosen $\left[11\right]$ to the quantum description of quantum nonlocality. Also this quantum description can be seen as a unitary evolution (the process of spatial separations of the two subsystems) between two one-to-one correlated measurement outcomes (those of measuring in the same basis the two respective maximally entangled observables) – we must think that this process goes from one to the other measurement outcome via the time the two subsystems were not spatially separate. Time symmetrized, it generates a quantum superposition of causal loops that play the role of the hidden variables envisaged by those authors to physically (locally) explain quantum non-locality $\left[5\right]$.
This ends the first part of the work, a review and further clarification of the retrocausal approach to the quantum speedup.
Of course, under the above explanation of their quantum speedup, quantum algorithms turn out to be teleological evolutions. The second part of the work is about the possibility that the quantum causal loops that stand at the basis of the teleological character of quantum algorithms are the missing physical basis of the teleological character of the evolutions of the living.
This immediately raises two problems.
One is explaining the presence of quantum features (the quantum speedup) in evolutions that can be the size of the Earth and have durations of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels points to the only possible solution: to scale them down from the largest and longest existing evolution there is, that of the universe. Of course, at the same time, we should adopt quantum cosmology, that is, the assumption that the universe – a perfectly isolated system since by definition there is nothing outside of it – evolves in a unitary quantum way.
The other is to find a formal analogy between quantum algorithms and, under quantum cosmology, natural evolutions. Obviously, it should justify the fact that also natural evolutions can occur with a quantum speedup and are consequently teleological in character.
Interestingly, the Anthropic Principle comes to our rescue. The Fine-tuned Universe version of it asserts that the unitary evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today, in particular (“sentient life” or “no life”). The reason for this correlation is that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Having adopted quantum cosmology, of course, also implies a quantum interpretation of the Anthropic Principle. Simply, the correlation between the value of the physical constants since the beginning of time and the value of life today becomes quantum and its building up is performed by the universe’s unitary quantum evolution.
This is what we were looking for. In fact, it is exactly what quantum algorithms do. They unitarily build up a quantum correlation between the problem-setting since the beginning of them and the solution of the problem at their end. Under this formal analogy, the evolution of the universe toward the value of life could be endowed with quantum speedup and would even in the case that the quantum search for the correlation in question is random, exactly as in random quantum search algorithms $\left[12\right]$.
Of course, being endowed with quantum speedup, under the quantum causal loops explanation of it, the evolution in question is teleological in character. In the universe we are in, what becomes teleological in character is the evolution of the universe toward sentient life.
Eventually, as we shall see, the evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out everything else, even if decohered, should inherit the quantum speed and thus the teleological character of the latter evolution. The same would consequently hold for any evolution of life on Earth, e. g. that of species.
The notion of a teleological evolution of life is of course puzzling. For example, the Darwinian theory of evolution would change as as follow: those random DNA mutations that, in the future, will survive natural selection would become more probable. With the past influenced by the future, this is a completely new way of viewing the physical world. No wonder the teleological notion has been rejected by science.
In conclusion, the work is highly unconventional, but its findings were obtained in a surprisingly simple and well-verifiable way. If it were correct, as we think, it would be a paradigm shift not only for the natural sciences, but also for the way we see physical reality.
@@ 2 Rationale
As the present work is about a fundamentally different way of viewing the role of causality in quantum (correlation) processes, before going to the formalizations we provide a Rationale that describes it conceptually.
This Rationale is divided into two parts corresponding to the two parts in which the work is divided.
The first provides the concept of the time-symmetric quantum causal loops, showing how they stand at the physical basis of the quantum computational speedup and quantum nonlocality. In particular, the time-symmetric quantum causal loop description of quantum algorithms highlights their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution of the problem again).
The second addresses the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
@@ 2.1 The time-symmetric quantum causal loops – the concept
We go now to the first part of the work. We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[6\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. Of course, we must ignore the ban on the use of the notion of retrocausality.
The fact that Alice (the problem-solver) always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
@@
We consider now $N$ drawers. Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
However, the above immediately raises a problem. The current quantum description of quantum algorithms, as that of any quantum process, is causal in character. In other words, causality always goes in the same time direction. Obviously, it is a description that cannot describe causal loops; we tentatively ascribe this to its incompleteness.
Before going to this assumed incompleteness, of a fundamental kind, we should emend a trivial incompleteness of the current quantum description of quantum algorithms. It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is obviously incomplete because it lacks the initial measurement. Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings. In the case of Grover algorithm, where the solution is an invertible function of the problem-setting (both are the number of the drawer with the ball), we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement). Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
By the way, seeing quantum algorithms as unitary evolutions between an initial and a final one-to-one correlated measurement outcomes is a key step of the present work. In view of going beyond Grover algorithm, we note that, if the solution is not an invertible function of the problem-setting, we can have one-to-one correlation between the initial and final measurement outcomes as well by requiring that, at the end, the problem-solver measures both the solution and the problem setting. The only constraint is that the quantum algorithm determines the solution with certainty.
We call the above, less incomplete, quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops. We should complete this ordinary quantum description in a way that leaves it mathematically unaltered but changes its causality structure – describes the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup. This is simply done time-symmetrizing it as follows.
@@
This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm. Note that, mathematically, i.e. disregarding the time-direction of causality, it is the ordinary quantum description back again.
Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure). As seen by the problem-solver, from whom the outcome of the initial measurement should always be concealed (would give her information about the solution – see Subsection 4.2.), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops. Of course, the backward in time propagations of all time-symmetrization instances are identical with each other. Therefore, as required, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered. However, it changes the structure of causality along it. As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; then, she can use this knowledge to produce the solution with fewer computation steps. If she exploits it entirely, the quantum speedup is optimal.
@@
In the present work, we will always know that there is such a unitary evolution. It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[4\right]$. Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics. Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops. Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup that, involving only the measurements of commuting observables, is submitted to the same (classical) logic. In turn, it is easily seen that the only way of obtaining these causal loops is by time-symmetrizing the ordinary quantum description of the quantum algorithm. In fact, to have a causal loop, we must properly share the selection of the pair problem-setting and corresponding solution between the initial and final measurements. Because of the time-symmetry of the unitary evolution between the two one-to-one correlated measurement outcomes, we should share it evenly. This exactly specifies the time-symmetrization we are dealing with. Therefore, as long as logic and quantum mechanics are consistent, the time-symmetric quantum causal loops generated by the time-symmetrization operation should be exempt from the paradoxes that plague the usual kind.
@@
In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered. When the two measurements become simultaneous, the time-symmetrization in question must already be there to become the atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
@@
There is also the case that the unitary evolution does not build up any quantum correlation. It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation. Here maximal entanglement between the two observables is already there before the space-separation process. After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate. Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[5\right]$. We can see that these causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e. also physically) describe quantum nonlocality.